An infant weights 37 kg. What is the required IV flow rate in mL/hr to maintain proper fluid levels?
Answer #1

For infant with a weight range of 20 to 70 kgs the required amount of fluid in a day is 1500 ml + 20 ml per each kg above 20 kgs

The weight of infant is 37 kgs = 37 - 20 = 17 kgs of additional fluid

Daily required fluid is 1500 ml + (20 ml / kg \times 17 kgs) = 1500 + 340 = 1840 ml / day

Flow rate = Volume in ml \div time in hours

                = 1840 \div 24

                = 76.6 ml

The infant requires to be maintained the intravenous fluid at the rate of 77 ml per hour.